OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x-y=-3 and passes through the point (4,-2)

OpenStudy (anonymous):

first we have to find the slope of required equation 3+3x=y slope is the coefficient of x. i.e.3 dis slope will be the slope of required line bcoz both the lines are parallel. given point is (4,2)=x,=4 and y1=2 (y-y1)=m(x-x1) put the values in dis equn u will gwt the ans